The current theory of the structure of the
Earth, called plate tectonics, tells us that the
continents are in constant motion.
Assume that the North American continent
can be represented by a slab of rock 4850 km on a side and 38 km deep and that the rock
has an average mass density of 2650 kg/m3.
The continent is moving at the rate of about
3.6 cm/year.
What is the mass of the continent?
Answer in units of kg
Don't round answer
1
Expert's answer
2019-11-04T10:11:39-0500
"m=\rho V=\rho a^2h"

"m=(2650)(4850000)^2(38000)=2.37\cdot 10^{21}kg"